How many grams of NH4Cl must be added to 0.100 L of 0.426 M NH3 to produce a buffer solution with pH= 9.34 ? pKa (NH4Cl) = 9.24 Units are not required on the answer. Round your answer to two decimal places.
Answer #1

pH = 9.34

pKa = 9.24

moles of NH3 = 0.100 x 0.426 = 0.0426 mol

pH = pKa + log [base / acid]

9.34 = 9.24 + log [0.0426 / NH4+]

[0.0426 / NH4+] = 1.26

moles of NH4+ = 0.0338 mol

moles of NH4Cl = 0.0338 mol

mass of NH4Cl = 0.0338 x 53.4915

mass of NH4Cl = 1.81 g

moles of NH4+ =